40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Aug 11 17:34:16 2017, awade, Craig, DailyProgress, BEAT, Mapping the beat notes north temp vs south temp NorthLaserTempVsPrecavityBeatnoteFrequency.pdf
    Reply  Sun Aug 13 19:58:49 2017, awade, Craig, DailyProgress, BEAT, Fixing the calibration of North laser Lightwave series 125/126 set point temperature 
       Reply  Mon Aug 14 16:34:45 2017, awade, Craig, DailyProgress, BEAT, Voltage offsets on unterminated slow frequency control on Lightwave lasers 2017-08-14_16.14.12.jpg2017-08-14_16.14.25.jpg
Message ID: 1858     Entry time: Mon Aug 14 16:34:45 2017     In reply to: 1857
Author: awade, Craig 
Type: DailyProgress 
Category: BEAT 
Subject: Voltage offsets on unterminated slow frequency control on Lightwave lasers 

Craig refound the BN this morning for the pre-cavity beat note detector.  Level is -41 dBm @ 360 MHz with northSlow = 39.7679 V and southSlow = 49.5736 (at the new laser controller calibration). This was with the EPICS slow controls unplugged at the front.  When I went to plug them back into the slow controls, both set to zero volts, the beat note went away.

After some trouble shooting I stuck a multimeter on the slow inputs and found that the slow inputs to the lasers' frequency control have a offset when unterminated.  North slow voltage offset is 3.151 V and south is -0.927 V (this is at the above temperature settings). These voltages change with the above setpoint temperatures unless the input is terminated or a voltage is applied.

The manual seems to suggest that the slow frequency BNC is an offset to the set point temperature. However, it seems like the BNC voltage input overrides the set point value rather than adding/subtracting from it.  If override is the case then full range of the laser frequency tuning should happen from -10 to +10 V on the slow inputs.  I will modify the relevant EPICS channels from -2 to 7 V to their full range of ±10 V.

---

Laser setpoint temperatures were reset back to the center of their range (48.0010 C South and 44.8010 C north).  The equivalent slow input voltages to reach the same 360 MHz BN is -1.2270 V south and 4.0870 V north. No more tuning on the laser controlers.  Epics channels from this point forward.

 

Attachment 1: 2017-08-14_16.14.12.jpg  2.172 MB  | Hide | Hide all
2017-08-14_16.14.12.jpg
Attachment 2: 2017-08-14_16.14.25.jpg  1.953 MB  | Hide | Hide all
2017-08-14_16.14.25.jpg
ELOG V3.1.3-